OpenStudy (anonymous):

find the area of the region bounded by the lines y=x and y=8x and the curve y=(1/x^2)

OpenStudy (mathmate):

The two lines intersect at the origin (0,0). The line y=8x intersects the curve y=1/x^2 at (1/2,4). The line y=x intersects the curve at (1,1). Will split the integrals into two, since the upper limit changes from the steep line to the curve. \[I1=\int\limits_{0}^{1/2}(8x-x)dx\] and \[I2 = \int\limits_{1/2}^{1} (1/x ^{2} - x)dx\] Evaluating the integrals, we get I1=7/8 and I2=5/8 for a total area of 3/2.